# Where Visual Speech Meets Language: VSP-LLM Framework for Efficient and Context-Aware Visual Speech Processing

### v1

@@ 4.3.2 Effectiveness of Rich Context Modeling
We have developed a unified model incorporating Large Language Models (LLMs) to leverage their advanced context modeling capabilities. Therefore, in this section, we conduct a qualitative experiment to demonstrate the effectiveness of the proposed VSP-LLM in handling homophenes, a challenging problem that requires substantial context understanding to accurately identify homophenes. Figure 3 shows several transcription examples obtained from AV-HuBERT and our model, illustrating how our proposed method accurately generates words by considering the entire context of a sentence. For instance, in a homophene case, AV-HuBERT incorrectly transcribes "barrow", a word that looks similar on the lips visually but differs in meaning from “marrow”. On the other hand, our method correctly generates "marrow", leading to complete the accurate phrase "bone marrow" in the context of describing a cancer treatment process. Similarly, AV-HuBERT’s transcription of "hair" is contextually inappropriate for a sentence discussing considerations regarding building systems. Our method, however, accurately outputs "air," resulting in the correct phrase "air system". Also, we can observe similar results in the other cases, not the homophene problem only. For example, the proposed method can generate the word “planetary” in the context of asking about the relationship between greenhouse gases and “planetary”. These results corroborate that our approach can more effectively comprehend contextual clues and generate more precise and natural answers, due to the integration of LLM.
Additionally, we evaluate the VSR performance across various video length segments to explore the effectiveness of LLM in handling long speech. Figure 4 shows that WER decreases as video length increases. Notably, our proposed method exhibits outstanding recognition performance, with a WER of 13.9% on videos longer than 6 seconds. Furthermore, our method demonstrates consistent performance improvements as the length of the video increases, compared to other methods. It indicates the effectiveness of LLM’s context modeling in longer video utterances, which demand a more comprehensive understanding of context.
  Number of Clusters BLEU normal-$\uparrow$ Length of sequence FLOPs (P)
@@ 4.3.3 Effectiveness of Deduplication
We conduct experiments to assess the effectiveness of our deduplication strategy. For the deduplication process, the number of clusters for visual speech units is required to be determined, and we show the effectiveness according to the number of clusters. Table 3 presents these results, and the first row shows the performance of the baseline which does not utilize the deduplication. The baseline obtains an average BLEU score of 22.0 with 27.2 peta FLOPs per training epoch. In contrast, our method acquires comparable performance, while significantly reducing the sequence length and computational resources (FLOPs). Specifically, with 200 clusters for visual speech units, our method not only maintains a similar performance level with a 22.1 average BLEU score but also cuts the sequence length by 53%. Consequently, the FLOPs are greatly reduced to 17.9, marking a 34.2% decrease. These experiments confirm that deduplication, applied to visual speech units, effectively eliminates redundant information.
@@ 4.3.4 VSP-LLM in Data-limited Situation
Leveraging the contextual understanding capabilities of LLM, which are pre-trained on vast text corpora, we suppose that a small amount of labeled data is sufficient for constructing a unified VSR and VST model. This is because the proposed VSP-LLM endeavors to establish the visual-to-text mapping while entrusting the task of language modeling to the LLM. To validate it, we train VSP-LLM with different amounts of labeled data; 1 hour, 5 hours, 10 hours, and 15 hours. For comparison, we also develop AV-HuBERT Shi et al. (2022) on the same data. Table 4 displays the VSR and VST performances. In all experimental conditions, regardless of the amount of data used, our proposed method significantly outperforms AV-HuBERT. Moreover, when using only 15 hours of labeled data, our unified method achieves a WER of 33.3%. This is a noteworthy achievement, particularly when compared to the previous VSR Makino et al. (2019) model achieving a WER of 33.6%, by using 31k hours of labeled data for training.
  Method Labeled Data(hrs) BLEU normal-$\uparrow$ WER(%) normal-$\downarrow$
@@ 5 Conclusion
In this paper, we proposed a novel framework, Visual Speech Processing with LLMs (VSP-LLM), designed to leverage the context modeling ability of LLMs. Through this framework, we built a unified model that can perform multi-tasks, VSR and VST, with a single model. Moreover, the proposed deduplication strategy reduces the redundant information of visual speech representations based on pronunciation information modeled from visual speech units. Through extensive experiments, we verified that the proposed deduplication method can reduce the visual sequence length by about 50% without dropping the performance. In addition, we validated the effectiveness of the VSP-LLM by achieving a superior performance in the MuAViC benchmark with only 15 hours of labeled data.
@@ Appendix A Visualization of Visual Speech Units
Appendix B Examples of Predicted Sentences
@@
Figure 6: Visualization of video frames corresponding to visual speech units. Each number indicates an index of visual speech unit.
Figure 7: Examples of VSR predictions produced by our proposed model on LRS3 test set. Deletions from the ground-truth text are highlighted in Red, while substitutions are shown in Blue.

### v2

@@ 4.3.2 Effectiveness of Rich Context Modeling
We have developed a unified model incorporating LLMs to leverage their advanced context modeling capabilities. Therefore, in this section, we conduct a qualitative experiment to demonstrate the effectiveness of the proposed VSP-LLM in handling homophenes, a challenging problem that requires substantial context understanding to accurately identify homophenes. Figure 3 shows several transcription examples obtained from AV-HuBERT and our model, illustrating how our proposed method accurately generates words by considering the entire context of a sentence. For instance, in a homophene case, AV-HuBERT incorrectly transcribes "i", a word which visually resembles "eye" on the lips, but differs in meaning. On the other hand, our method correctly generates "eye", successfully completing the idiom "eye to eye" to describe mutual understanding between individuals. Similarly, AV-HuBERT’s transcription of "write" is contextually inappropriate for a sentence discussing teaching the physical skill of riding a bike. Our method, however, accurately outputs "ride" resulting in the correct phrase "ride a bike". Also, we can observe similar results in the other cases, not the homophene problem only. For example, the proposed method can generate the word “composite” according to standard English usage, unlike AV-HuBERT, which erroneously outputs "compositive". These results corroborate that our approach can more effectively comprehend contextual clues and generate more precise and natural answers, due to the integration of LLM.
Additionally, we evaluate the VSR performance across various video length segments to explore the effectiveness of LLM in handling long speech. Figure 4 shows that WER decreases as video length increases. Notably, our proposed method exhibits outstanding recognition performance, with a WER of 12.9% on videos longer than 6 seconds. Furthermore, our method demonstrates consistent performance improvements as the length of the video increases, compared to other methods. It indicates the effectiveness of LLM’s context modeling in longer video utterances, which demand a more comprehensive understanding of context.
  Number of Clusters BLEU $\uparrow$ Length of sequence FLOPs (P)
@@ 4.3.3 Effectiveness of Deduplication
We conduct experiments to assess the effectiveness of our deduplication strategy. For the deduplication process, the number of clusters for visual speech units is required to be determined, and we show the effectiveness according to the number of clusters. Table 3 presents these results, and the first row shows the performance of the baseline which does not utilize the deduplication. The baseline obtains an average BLEU score of 14.6 with 62.4 peta FLOPs per training epoch. By applying the proposed deduplication, our method acquires comparable performance, while significantly reducing the sequence length and computational resources (FLOPs). Specifically, with 200 clusters for visual speech units, our method not only maintains a similar performance level with a 14.5 average BLEU score but also cuts the sequence length by 53%. Consequently, the FLOPs are greatly reduced to 45.6, marking a 26.9% decrease. These experiments confirm that deduplication, applied to visual speech units, effectively eliminates redundant information.
@@ 4.3.4 VSP-LLM in Data-limited Situation
Leveraging the contextual understanding capabilities of LLM, which are pre-trained on vast text corpora, we suppose that a small amount of labeled data is sufficient for constructing a unified VSR and VST model. This is because the proposed VSP-LLM endeavors to establish visual-to-text mapping while entrusting the task of language modeling to the LLM. To validate it, we train VSP-LLM on the MuAViC dataset with different amounts of labeled data; 1 hour, 5 hours, 10 hours, and 15 hours. For comparison, we also develop AV-HuBERT on the same data. Table 4 displays the VSR and VST performances. In all experimental conditions, regardless of the amount of data used, our proposed method significantly outperforms AV-HuBERT. Moreover, when using only 15 hours of labeled data, our unified method achieves a WER of 32.8%. This is a noteworthy achievement, particularly when compared to the previous VSR Makino et al. (2019) model achieving a WER of 33.6%, by using 31k hours of labeled data for training.
  Method Labeled Data(hrs) BLEU $\uparrow$ WER(%) $\downarrow$
@@ 5 Conclusion
In this paper, we proposed a novel framework, Visual Speech Processing with LLMs (VSP-LLM), designed to leverage the context modeling ability of LLMs. Through this framework, we built a unified model that can perform multi-tasks, VSR, and VST, with a single model. Moreover, the proposed deduplication strategy reduces the redundant information of visual speech representations based on pronunciation information modeled from visual speech units. Through extensive experiments, we verified that the proposed deduplication method can reduce the visual sequence length by about 50% with minimal performance degradation. In addition, we validated the effectiveness of the VSP-LLM by achieving a superior performance in the MuAViC benchmark with only 30 hours of labeled data.
6 Limitations
We have proposed a powerful visual speech processing method that incorporates LLMs to recognize and translate lip movements into other languages, leveraging the rich context modeling ability of LLMs. Despite the impressive improvement in the performance of this proposed method, the utilization of LLMs has been limited to VSR and VST tasks. We expect that the proposed VSP-LLM framework can be expanded to in real-world communication scenarios by utilizing additional non-verbal cues such as facial expressions and gestures. Especially, the VSP-LLM combined with non-verbal cues is expected to perform various tasks such as emotional recognition and dialog generation, starting with this paper as a foundation.
7 Broader impact and ethics
The integration of Large Language Models (LLMs) within our framework plays a pivotal role in its ability to handle the complexities of visual speech across different languages. LLM brings a deep understanding of contextual and linguistic information, which is critical for accurately interpreting and translating visual speech cues. This capacity for nuanced language processing underpins our confidence in the framework’s potential for broader linguistic applicability. Moreover, our experiments have demonstrated exceptional data efficiency and significant performance gains with relatively small amounts of labeled data for each language. This efficiency is crucial for scalability to other languages and dialects, particularly those for which extensive labeled datasets may not be readily available. The ability to achieve robust performance with limited data is indicative of the framework’s adaptability and its potential for expansion to a wider linguistic range.
@@ Appendix A Visualization of Visual Speech Units
Appendix B FLOPs During Inference with Deduplication
Table 5 shows the FLOPs during inference time. Similar to during training, applying deduplication techniques also significantly reduced inference FLOPs.
  Sample ID Label
test/VIgzTLDyObo/00004 and then what happens
trainval/jpeSLKnS4gM/50020 and then what happens
[…]
pretrain/omGbKQIzoWY/00009_00 you probably won’t do well on that problem on the other hand relaxed daydreaming is a way to
Table 6: Examples of cases where sentences in the test set also appear in the training set, but are spoken by distinct individuals.
Appendix C Exposure to Transcriptions in the Pre-Training of LLM
There might be concerns regarding LLaMA2’s potential exposure to the LRS3 dataset during the pre-training phase. Since the details of LLaMA2’s training data aren’t publicly available, we can’t be absolutely sure whether LRS3 was included or not. However, it’s important to emphasize that the core challenge and focus of visual speech recognition (VSR) and translation (VST) lie in the ability to accurately match mouth shapes to unseen speakers, rather than merely replicating text from specific sentences. In particular, the mouth shape of the same sentence can vary significantly when expressed by different speakers, emphasizing the visual rather than textual nature of the work. Our analysis of the LRS3 dataset (Table 6) highlights this point, showing cases where sentences in the test set also appear in the training set, but are spoken by distinct individuals. This case serves to highlight the importance of the model’s ability to recognize speaker-specific mouth shapes over memorizing textual content. Given this context, we believe that the potential exposure of LLaMA2 to certain sentences from the LRS3 dataset during training is unlikely to significantly impact the model’s performance in our study.
Homophene Cases
Ground Truth it’s not like teaching them how to ride a bike
Prajwal et al. (2022) it’s all i teach them how to write a bike
VSP-LLM it’s not like teaching them how to ride a bike
Ground Truth is it about earning as much as you possibly can
Prajwal et al. (2022) it’s about learning as much as possibly can
VSP-LLM it’s about earning as much as you possibly can
Ground Truth it’s like a piece of junk mail to be thrown away
Ma et al. (2021b) it’s like a piece of chunk made to be thrown away
VSP-LLM it’s like a piece of junk mail being thrown away
Ground Truth and imagine what might happen because every region has something to offer
Ma et al. (2021b) and imagine what might happen because every reason has something to offer
VSP-LLM and imagine what might happen because every region has something to offer
Table 7: Additional baseline examples for the homophene case. The Red words indicate homophene words.
Appendix D Additional Examples of Homophene case
In Section 4.3.2, we discussed the VSP-LLM model’s exceptional ability to correctly distinguish homophenes by leveraging its advanced context modeling capabilities. This section further extends our analysis by comparing the performance of the VSP-LLM with other baseline models in handling homophenes. The results of these comparisons are presented in Table 7. In one notable example, Ma et al. incorrectly transcribed "junk" as "chunk." In contrast, the VSP-LLM accurately recognized the phrase "junk mail," a commonly used and contextually appropriate phrase in English. This illustrates the VSP-LLM’s superior performance, particularly its proficiency in integrating contextual understanding with linguistic patterns to enhance transcription accuracy in cases involving homophenes.
Figure 7: Examples of VSR and VST predictions produced by our proposed model on LRS3 and En-to-X test set. Deletions from the ground-truth text are highlighted in Red, while substitutions or addition are shown in Blue.
Appendix E Examples of Predicted Sentences
@@
Appendix F Erratum
In Section 4 of this paper, we intended to report the translation results on the MuAViC dataset. However, the results were mistakenly reported on the LRS3-T dataset instead of the MuAViC dataset. In the revised manuscript, we have corrected this error by assessing and reporting the translation results on the correct MuAViC dataset.